an executive invents $21,000, some at 7% and the rest at 4% annual interests. if he receives an annual return of $1,320, how much is invested at each rate?
Answer by addingup(3677)

0.07x + 0.04(21,000-x) = 1320
0.07x + 840-0.04x = 1320
0.03x = 480
x = 16,000 this is the amount at 7%
The rest: 21,000-16,000 = 5,000 this is the amount at 4%